# Robust Design and Evaluation of

### v5

@@ Appendix E Additional Tables for Monte Carlo Simulations and Empirical Application
Table 2: Bias and coverage properties with nonparametric outcome bounds.
[…]
Notes: This table summarizes the average bias of $\widehat{\overline{\mbox{perf}}}(s;\beta)$ for the upper bound on MSE and $\widehat{\underline{\mbox{perf}}}(s;\beta)$ for the lower bound on accuracy, the standard deviations of our estimators (SD), their average estimated standard errors ($\widehat{\sigma}$), and the coverage rate of nominal 95% confidence intervals based on the asymptotic normal approximation in Proposition 4.2. We report these results for $n\in\{500,1000,2500\}$ and nonparametric outcome bounds with $\underline{\Gamma}=2/3$, $\overline{\Gamma}=3/2$. The results are computed over 1,000 simulations. See Section 5 for further discussion.
@@
Table 3: Detailed description of variable names in right panel of Table 3b.

### v6

@@ F.2 Behavior of Estimated Predictive Performance Bounds
Table A2: Bias and coverage properties with observed outcome bounds.
(a) Upper bound on mean square error
(b) Lower bound on accuracy
Figure A4: Distribution of overall performance estimators across Monte Carlo simulations with observed outcome bounds.
Appendix G Additional Tables for Empirical Application
In this section, we report an additional table for the empirical application that is referenced in Section 5 of the main text. Table A3 provides detailed descriptions of the variable names in right panel of Figure 1 in Section 5.1 of the main text. Personal loans applications can have multiple listed applicants, so some variables refer to just the first listed applicant.
@@ Appendix G Additional Tables for Empirical Application
Table A3: Detailed description of variable names in right panel of Table 1.